OpenStudy (anonymous):

what is the solution to the equation 2/5 x=20 A. 10 B. 50 C. 40 D. 8

OpenStudy (anonymous):

is it like this:\[\frac{ 2 }{ 5 }x = 20\]

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

there is a fraction there , and we want to get rid of the fraction, do you know how we can best do that? Removing the fraction part?

OpenStudy (anonymous):

@skjdbgbsbngvolb

OpenStudy (anonymous):

ill give medal and be your faan

OpenStudy (anonymous):

ok, you get rid of the denominator by multiplying both sides by 5(the denominator), such that:\[5 \times \frac{ 2 }{ 5 }x = 20 \times5\] \[2x = 100\] \[\frac{ 2 }{ 2 }x = \frac{ 100 }{ 2 }\] \[x= 50\]